OpenStudy (anonymous):

Gregory has a coupon for $1 off the purchase of 3 boxes of Best Brand cereal. The store has 5 varieties of Best Brand cereal. How many ways can Gregory choose 3 different types of cereal?

OpenStudy (nurali):

combination formula. n = 5 and r = 3, so the number of ways is 5!/((5-3)!3!) = 120/(2·6) = 10.

OpenStudy (anonymous):

thank you!!!!!!